OpenStudy (anonymous):

\[|x-y| + |y-z| = |x - y + y-Z|\] I do not understand the steps between the left side of the equation and the right side of the equation. Please explain.

OpenStudy (jamesj):

This relation is false. Let x = 0, y = 1 and z = 0 then |x-y| + |y-z| = 1 + 1 = 2 but |x - y + y - z| = |x - z| = 0

OpenStudy (anonymous):

I wrote down the problem incorrectly.